OpenStudy (anonymous):

Let f(n) = n2 + 2n + 4 and g(n) = - 3n + 2 Find (f o g)(1)Answer

OpenStudy (anonymous):

f(n)=n^2+2n+4 replace n with g(n) (f o g)(n)=(-3n+2)^2+2(-3n+2)+4 now evaluate in 1: (fog)(1)=(-3(1)+2)^2+2(-3(1)+2)+4= 3